### v1

@@ Abstract
Methods:
We employed the nnU-Net framework to evaluate eleven annotation strategies for accurate 3D segmentation of pheochromocytoma, introducing a set of novel multi-class schemes based on organ-specific anatomical priors. These priors were derived from adjacent organs commonly surrounding adrenal tumors (e.g., liver, spleen, kidney, aorta, adrenal gland, and pancreas), and were compared against a broad body-region prior used in previous work. The framework was trained and tested on 105 portal venous phase contrast-enhanced CT scans from 91 patients in a retrospective cohort imaged at the NIH Clinical Center. Performance was assessed using Dice Similarity Coefficient (DSC), Normalized Surface Distance (NSD), and instance-wise F1 score.
@@ 1 Introduction
Pheochromocytoma (PCC) is a rare neuroendocrine tumor arising primarily from the chromaffin cells of the adrenal medulla [1]. Characterized by excessive secretion of catecholamines, pheochromocytomas can lead to significant cardiovascular morbidity, including severe hypertension, palpitations, and potentially life-threatening hypertensive crises [2]. Besides, about 10-15% of PCCs are metastatic [3, 4]. These tumors are unique in their highly variable molecular landscape driven by genetic alterations, either germline or somatic [5]. These mutations translate into three major genetic clusters: pseudohypoxia-related cluster 1A-SDHx and cluster 1B-VHL/EPAS1 mutations, cluster 2-kinase signaling (such as RET, NF1, TMEM127 and MAX), and a less defined group Wnt signaling–related cluster 3, often comprising sporadic cases [6, 7, 8]. The underlying genetic cluster can influence tumor location, biochemical profile, metastatic potential, and overall clinical management strategy [9, 10]. Timely and accurate diagnosis, localization, and characterization of these tumors are crucial for optimal management, including surgical resection, which remains the primary therapeutic intervention [11, 12]. Computed tomography (CT), especially contrast-enhanced abdominal imaging, is routinely employed due to its high spatial resolution, wide availability, and rapid imaging capability, making it the imaging modality of choice for initial detection and characterization of adrenal masses [13]. Figure 1 illustrates the portal venous phase CT volumes containing PCC tumors from different genetic clusters. A 2D slice from each volume are displayed with a soft tissue window center and width of [50, 450] HU.
@@
Recent advances in artificial intelligence (AI), particularly deep learning-based segmentation approaches such as U-Net and its variants, have demonstrated substantial improvements in medical image segmentation tasks, including tumor delineation [20, 21, 22]. Despite these advancements, the segmentation of pheochromocytomas remains challenging due to variability in tumor size, shape, contrast uptake patterns, and proximity to adjacent anatomical structures, making it difficult for traditional and even some contemporary automated methods to achieve consistently high accuracy. Among recent AI-based segmentation methods, the nnU-Net segmentation method is the most widely adopted in various biomedical segmentation applications because of its high accuracy and efficiency [23]. While the literature on pheochromocytoma segmentation from abodminal CT is limited, some prior approaches proposed training nnU-Net exploiting 2D or pseudo-3D annotations [24, 25]. Additionally, those approaches addressed PCC detection by predicting bounding boxes, rather than precise tumor delineation or segmentation. Such detection-based 2D/pseudo-3D models are not suitable for the accurate quantification of key tumor characteristics, such as volume, boundary irregularity, and changes in standardized uptake values (SUV), which are crucial for assessing tumor progression.
@@
Figure 1: The portal venous phase CT scans of Pheochromocytoma tumors from different genetic subtypes. Red contours indicate the tumor locations in each scan.
@@ 2.4 Experimental Settings
Training Parameters. Training was performed using the nnU-Net v2.5.2 framework with a 3D full-resolution configuration. The model was trained with a batch size of 2 and a patch size of 80$\times$ 160$\times$ 160, using CT volumes resampled to a uniform spacing of [2.5, 0.774, 0.774] mm and normalized with default CT intensity normalization. The architecture followed a 6-stage 3D U-Net (PlainConvUNet) with increasing feature maps from 32 to 320, instance normalization, and LeakyReLU activation. Optimization was performed using stochastic gradient descent (SGD) with an initial learning rate of 0.01, momentum of 0.99, and weight decay of 3 ×$\times 10^{-5}$. A polynomial learning rate scheduler was used to gradually reduce the learning rate across 1000 training epochs. The model was trained using a NVIDIA A100 GPU in the NIH High Performance Computing (HPC) cluster.
@@ 4 Discussion and Conclusion
In conclusion, the findings of our study highlight the critical role of carefully selecting anatomical priors in deep learning-based segmentation tasks. The proposed segmentation framework represents a substantial advancement in automated pheochromocytoma segmentation, offering the potential to improve clinical workflows, enable more reliable disease monitoring, and support enhanced prognostic assessments. Moreover, accurate PCC segmentation may serve as a valuable tool for predicting the underlying genetic cluster associated with the tumor, providing a cost-effective alternative to genetic testing for initial risk stratification. Future work will investigate integrating these segmentation models into broader clinical decision-support systems and evaluating their generalizability across larger, multi-institutional datasets. Additionally, the framework could be adapted for segmenting paragangliomas (non-adrenal tumors), potentially incorporating additional imaging context, such as functional PET imaging, to better localize tumors that are often small, heterogeneous, and difficult to identify on CT alone.

### v2

@@ Abstract
Methods: We employed the nnU-Net framework to evaluate eleven annotation strategies for accurate 3D segmentation of pheochromocytoma, introducing a set of novel multi-class schemes based on organ-specific anatomical priors. These priors were derived from adjacent organs commonly surrounding adrenal tumors (e.g., liver, spleen, kidney, aorta, adrenal gland, and pancreas), and were compared against a broad body-region prior used in previous work. The framework was trained and tested on 105 portal venous phase contrast-enhanced CT scans from 91 patients in a retrospective cohort imaged at the NIH Clinical Center. Performance was assessed using Dice Similarity Coefficient (DSC), Normalized Surface Distance (NSD), and instance-wise F1 score.
@@ 1 Introduction
Pheochromocytoma (PCC) is a rare neuroendocrine tumor arising primarily from the chromaffin cells of the adrenal medulla [1]. Characterized by excessive secretion of catecholamines, pheochromocytomas can lead to significant cardiovascular morbidity, including severe hypertension, palpitations, and potentially life-threatening hypertensive crises [2]. Besides, about 10-15% of PCCs are metastatic [3, 4]. These tumors are unique in their highly variable molecular landscape driven by genetic alterations, either germline or somatic [5]. These mutations translate into three major genetic clusters: pseudohypoxia-related cluster 1A-SDHx and cluster 1B-VHL/EPAS1 mutations, cluster 2-kinase signaling (such as RET, NF1, TMEM127 and MAX), and a less defined group Wnt signaling–related cluster 3, often comprising sporadic cases [6, 7, 8]. The underlying genetic cluster can influence tumor location, biochemical profile, metastatic potential, and overall clinical management strategy [9, 10]. Timely and accurate diagnosis, localization, and characterization of these tumors are crucial for optimal management, including surgical resection, which remains the primary therapeutic intervention [11, 12]. Computed tomography (CT), especially contrast-enhanced abdominal imaging, is routinely employed due to its high spatial resolution, wide availability, and rapid imaging capability, making it the imaging modality of choice for initial detection and characterization of adrenal masses [13]. Figure 1 illustrates the portal venous phase CT volumes containing PCC tumors from different genetic clusters. A 2D slice from each volume is displayed with a soft tissue window center and width of [50, 450] HU.
@@
Recent advances in artificial intelligence (AI), particularly deep learning-based segmentation approaches such as U-Net and its variants, have demonstrated substantial improvements in medical image segmentation tasks, including tumor delineation [20, 21, 22]. Despite these advancements, the segmentation of pheochromocytomas remains challenging due to variability in tumor size, shape, contrast uptake patterns, and proximity to adjacent anatomical structures, making it difficult for traditional and even some contemporary automated methods to achieve consistently high accuracy. Among recent AI-based segmentation methods, the nnU-Net segmentation method is the most widely adopted in various biomedical segmentation applications because of its high accuracy and efficiency [23]. While the literature on pheochromocytoma segmentation from abdominal CT is limited, some prior approaches proposed training nnU-Net exploiting 2D or pseudo-3D annotations [24, 25]. Additionally, those approaches addressed PCC detection by predicting bounding boxes, rather than precise tumor delineation or segmentation. Such detection-based 2D/pseudo-3D models are not suitable for the accurate quantification of key tumor characteristics, such as volume, boundary irregularity, and changes in standardized uptake values (SUV), which are crucial for assessing tumor progression.
@@
Figure 1: Portal venous phase abdominal CT scans showing pheochromocytoma tumors from different genetic subtypes. Red contours indicate the tumor locations in each scan. (b) and (c) show bilateral tumors.
@@ 2.4 Experimental Settings
Training Parameters. Training was performed using the nnU-Net v2.5.2 framework with a 3D full-resolution configuration. The model was trained with a batch size of 2 and a patch size of 80$\times$ 160$\times$ 160, using CT volumes resampled to a uniform spacing of [2.5, 0.774, 0.774] mm and normalized with default CT intensity normalization. The architecture followed a 6-stage 3D U-Net (PlainConvUNet) with increasing feature maps from 32 to 320, instance normalization, and LeakyReLU activation. Optimization was performed using stochastic gradient descent (SGD) with an initial learning rate of 0.01, momentum of 0.99, and weight decay of 3 ×$\times 10^{-5}$. A polynomial learning rate scheduler was used to gradually reduce the learning rate across 1000 training epochs. The model was trained using an NVIDIA A100 GPU in the NIH High Performance Computing (HPC) cluster.
@@ 4 Discussion and Conclusion
In conclusion, the findings of our study highlight the critical role of carefully selecting anatomical priors in deep learning-based segmentation tasks. The proposed segmentation framework represents a substantial advancement in automated pheochromocytoma segmentation, offering the potential to improve clinical workflows, enable more reliable disease monitoring, and support enhanced prognostic assessments. Moreover, accurate PCC segmentation can serve as a valuable tool for predicting the underlying genetic cluster associated with the tumor, providing a cost-effective alternative to genetic testing for initial risk stratification.
Despite the promising results of this study, certain limitations should be acknowledged. The dataset was retrospectively collected from a single institution, which may limit the generalizability of the findings to broader populations or imaging settings. Although five-fold cross-validation was performed to assess robustness, external validation using multi-center datasets with greater demographic and scanner variability would further support the clinical applicability of the proposed approach. Future work will focus on evaluating the generalizability of the proposed segmentation framework across larger, multi-institutional datasets and exploring its integration into broader clinical decision-support systems. Additionally, the framework could be extended to segment paragangliomas (non-adrenal neuroendocrine tumors), potentially incorporating additional imaging context, such as functional PET imaging, to better localize tumors that are often small, heterogeneous, and difficult to identify on CT alone.